OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to 2x+y=8 and passes through (0, 3)

OpenStudy (aroub):

same as that one so... y=mx+b y=-2x+8 3=-2(0)+b b=3 y=-2x+3

OpenStudy (anonymous):

YOU ARE THE BESTEST (: (: i really appreciate it

OpenStudy (aroub):

thank you :):)

OpenStudy (anonymous):

welcome (: